In an article in the Journal of Advertising, Weinberger and Spotts compare the use of humor in television ads in the United States and in the United Kingdom. Suppose that independent random samples of television ads are taken in the two countries. A random sample of 400 television ads in the United Kingdom reveals that 142 use humor, while a random sample of 500 television ads in the United States reveals that 124 use humor.

(a) Set up the null and alternative hypotheses needed to determine whether the proportion of ads using humor in the United Kingdom differs from the proportion of ads using humor in the United States.

H0: p1p2 (Click to select)=≠ 0 versus Ha: p1p2 (Click to select)=≠ 0.

(b) Test the hypotheses you set up in part a by using critical values and by setting α equal to .10, .05, .01, and .001. How much evidence is there that the proportions of U.K. and U.S. ads using humor are different? (Round the proportion values to 3 decimal places. Round your answer to 2 decimal places.)

  z = _________

(Click to select)RejectDo not Reject H0 at each value of α; (Click to select)nonesomestrongvery strongextremely strong evidence.

(c) Set up the hypotheses needed to attempt to establish that the difference between the proportions of U.K. and U.S. ads using humor is more than .05 (five percentage points). Test these hypotheses by using a p-value and by setting α equal to .10, .05, .01, and .001. How much evidence is there that the difference between the proportions exceeds .05? (Round the proportion values to 3 decimal places. Round your z value to 2 decimal places and p-value to 4 decimal places.)

z =
p-value =

(Click to select)Do not RejectReject H0 at each value of α = .10 and α = .05; (Click to select)somenoneextremely strongstrongvery strong evidence.

(d) Calculate a 95 percent confidence interval for the difference between the proportion of U.K. ads using humor and the proportion of U.S. ads using humor. Interpret this interval. Can we be 95 percent confident that the proportion of U.K. ads using humor is greater than the proportion of U.S. ads using humor? (Round the proportion values to 3 decimal places. Round your answers to 4 decimal places.)

95% of Confidence Interval                      [ , ]

(Click to select)YesNo the entire interval is above zero.
Answer #1

(a)

The null and alternative hypothesis are,

Ho: p1 - p2 = 0

HA: P1 - p270

Where, p1 is proportion of television ads uses humor in UK and p2 is proportion of television ads uses humor in USA

(b)

From the data values given,

142 P1 = 20 400 = 0.355

124 500 = 0.248

The z-test is used to compare two population proportion.

The z-statistic is,

P1 P2 Vo(1 – Đ) (+ )

Where,

= 21 +22 ny + n2 142 +124 4 = 0.2956 400+ 500

- 10.355 -0.248 ¥0.2956(1-0.2956) ( + : -3.4957 )

= 3.4957x3.50

Now, the critical values are obtained from standard normal distribution table for significance level = .10, .05, .01, and .001.

This is a right tailed test.

The critical values are,

For a=0.10 21-a = 21-0.10 = 1.28

Alpha Z-statistic Z-critical value
0.1 3.496 >    1.282 Significant, Reject the null hypothesis
0.05 3.496 > 1.645 Significant, Reject the null hypothesis
0.01 3.496 > 2.326 Significant, Reject the null hypothesis
0.001 3.496 > 3.090 Significant, Reject the null hypothesis

(c)

The null and alternative hypothesis are,

Ho: P1 - P2 < 0.05

HA: P1 - p2 > 0.05

The z-test is used to compare two population proportion.

The z-statistic is,

z=\frac{(p_1-p_2)-0.05}{\sqrt{\overline{p}(1-\overline{p})\left ( \frac{1}{n_1}+\frac{1}{n_2} \right )}}

Where,

z=\frac{(0.355-0.248)-0.05}{\sqrt{0.2956(1-0.2956)\left ( \frac{1}{400}+\frac{1}{500} \right )}}=1.8622

2= 1.8622 1.86

Now, the P- value is are obtained from z distribution table for Z-statistic = 1.8622.

P-value = 0.0313

P-value Alpha
0.0313 < 0.1 Significant
0.0313 < 0.05 Significant
0.0313 > 0.01 Not Significant
0.0313 > 0.001 Not Significant

At most 5% significance level, the difference between the proportions exceeds .05.

d)

The 95% Confidence Interval of the difference of two proportions is obtained by,

\text{95\% CI}=(p_1-p_2)\pm z\times\sqrt{ \frac{p_1(1-p_1)}{n_1}+\frac{p_2(1-p_2)}{n_2}}

\text{95\% CI}=(0.355-0.248)\pm 1.96\times\sqrt{ \frac{0.355(1-0.355)}{400}+\frac{0.248(1-0.248)}{500}}

\text{95\% CI}=\left ( 0.0467,0.1673 \right )

Entire interval is above zero which indicate that, at 5% significance level, the difference in mean is not zero.